OpenStudy (anonymous):

Tanya bought flowers for her mother she spent $45. Each tulip was $1 and each rose was $2 she bought 3 more Tulips than roses. let T be the number for tulips. Form an algebraic sentence to help you figure out how many tulips Tonya bought.

OpenStudy (anonymous):

Would it be 2 (t+3) = 45?

OpenStudy (zzr0ck3r):

t = r+3 because there are 3 more tulips than roses t+2r=45 because 1$*amount of tulips +2$*amount of roses = 45

OpenStudy (anonymous):

2(t+3)=45 2t + 6 = 45 t + 6 - 22.5 t=16.5 This implies that she purchased 16 and a half tulips...

OpenStudy (zzr0ck3r):

r+3+2r=45 3r+3=45 3r=42 r=14 t = 17

OpenStudy (anonymous):

The choices I have are. t + 2 (t+3) =45 t + 2 (t-3) =45 2t (t+3) =45 -3t + 2t =45

OpenStudy (texaschic101):

first equation is supposed to be 1t + 2r = 45 and r = t - 3 so I subbed and got : t + 2(t - 3) = 45

OpenStudy (zzr0ck3r):

t + 2 (t-3) =45 implies t = 17 and this is what I got, and I have faith in that.

OpenStudy (texaschic101):

I agree...thats what I got the second time

OpenStudy (zzr0ck3r):

17*1+14*2=17+28=45

OpenStudy (texaschic101):

the first time I subbed r + 3 in for t 1t + 2r = 45 and got: r + 3 + 2r =45